OpenStudy (anonymous):

2 tan^2x+11 tanx+9=0 plesssssss I have quizzzz

OpenStudy (anonymous):

help me plesssssss

OpenStudy (anonymous):

:(

OpenStudy (ddcamp):

Where are the parentheses on the equation?

OpenStudy (anonymous):

(0,2pi)

OpenStudy (anonymous):

replace tan(x) with x to make it easier and then you have 2x^2 +11x+9=0

OpenStudy (anonymous):

so what is the answer im not that good on this subject

OpenStudy (anonymous):

they need the exact value and approximate value

OpenStudy (anonymous):

well then you can factor it into (2x+9)(x+1) and solve each portion for x, replace x with tan(x) and take the inverse tangent

OpenStudy (anonymous):

Wilfley :( I don't know how to do it

OpenStudy (anonymous):

2x+9=0 x=-9/2 tan(x)=-9/2 tan^-1(-9/2)=x x+1=0 x=-1 tan(x)=-1 tan^-1(-1)=x

OpenStudy (anonymous):

then for approximate values of x just plug in to your calculator

OpenStudy (anonymous):

nice, now tan^-1(-1)=x this one is the exact value for x ?

OpenStudy (anonymous):

one of them, the other is the one above it

OpenStudy (anonymous):

choose it for me plessss :*

OpenStudy (anonymous):

\[x=\tan^{-1} (-\frac{ 9 }{ 2 }) and x=\tan^{-1} (-1)\]

OpenStudy (anonymous):

it should be between (0,2pi)

OpenStudy (anonymous):

Wilfleyyyyyyyyyyyy

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

trying to remember

OpenStudy (anonymous):

plessssssssss

OpenStudy (anonymous):

I think I'm missing something or just can't remember, for the -1 the exact is -pi/4 which I believe would just be moving clockwise from 0 pi/4 which would be between o and 2pi

OpenStudy (anonymous):

-1 is 3pi/4 I think

OpenStudy (anonymous):

I don't want to be wrong though

OpenStudy (anonymous):

yeah, -pi/4 is 3pi/4

OpenStudy (anonymous):

u know it more than me :(

OpenStudy (anonymous):

just a minute and I'll check it

OpenStudy (anonymous):

ples my life in ur hand lol

OpenStudy (anonymous):

you have to pick 1 of them?

OpenStudy (anonymous):

x+1=0